OpenStudy (anonymous):

medal and i'll fan everyone who answers. the sum of the digits of a two digit number is 12. the number formed by interchanging the digits is 54 more than the original number. what is the original number? A. 39 B. 48 C. 57

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

first equation says \(x+y=12\) if the number is \(xy\) second says \(10y+x=10x+y+54\)

OpenStudy (anonymous):

or if you prefer \[9y-9x=54\] or even \[y-x=6\]

OpenStudy (anonymous):

oops i meant \[x+y=12\\ y-x=6\]

OpenStudy (anonymous):

wow that is alot of numbers and letters. what do i do with the equations?

OpenStudy (anonymous):

you get \(2y=18\) so \(y=9\) making \(x=3\)

OpenStudy (anonymous):

the original number is \(39\) if you reverse the digits you get \(93\) which is \(39+54\)

OpenStudy (anonymous):

thank you so much you are so helpful. i'm sorry i'm really bad at this subject

OpenStudy (anonymous):

course since you had choices it would have been easier just to check reverse the digits of 39 (your first option) get 93 which is \(39+54\) done

OpenStudy (anonymous):

can you try and teach me how to do the next one if i type it on here?

OpenStudy (anonymous):

and ohh thats what we did the last time

OpenStudy (anonymous):

yw, i am sure you are not bad, just not in practice yet sure go ahead and ask, i will help if i can

OpenStudy (anonymous):

yeah, i didn't see the choices before solving

OpenStudy (anonymous):

the length of a rectangle is 2 more than 3 times the width. if the perimeter is 100, find the lenth and width of the rectangle. if l= length and w= width, which of the following systems could be used to solve the problem? A. l=3w+2 and 2l+2w=100 B. l=2w+3 and 2l+2w=100 C. l+2=3w and 2l+2w=100

OpenStudy (anonymous):

length is two more than three times the width \[l=3w+2\]

OpenStudy (anonymous):

so your choice is A because all of them have \(2l+2w=100\) as a second option

OpenStudy (anonymous):

okay I understand that thank you!! the next one is the length of a rectangle is 2 more than 3 times the width . if the perimeter is 100 meters what is width of the rectangle? A. 13 meters B. 12 meters C. 11 meters thanks for helping this is my last quiz of the semester and i have to get an a to have a on this quiz to bring my grade up to a decent grade to finish the semester with

OpenStudy (anonymous):

this one we check

OpenStudy (anonymous):

like a guess and check table?

OpenStudy (anonymous):

if the width is \(13\) then the length is \(3\times 13+2=41\) and the perimeter is \[2\times 13+2\times 41=108\] too big

OpenStudy (anonymous):

12 will work because \(3+\times 12+2=38\) and \[2\times 12+2\times 38=100\]

OpenStudy (anonymous):

typo there, i meant \[3\times 12+2=38\] but the answer is correct , it is 12

OpenStudy (anonymous):

and 2x11+2= 24 so you're right it has to be 12

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

do you mind helping me a few more?

OpenStudy (anonymous):

gotta run soon we can do one more or so

OpenStudy (anonymous):

dad is four times as old as his son jim. in ten years, dad's age will be 20 years more than twice jims age. how old is jim? if d= dads age now and j=jims age now, which of the following systems could be used to solve the problem? A.D=4J and D+10=2J+20 B. 4J and D +10=2(4J+10)+20 C. 4J and D +10=2(J+10)+20

OpenStudy (anonymous):

and how old is jim currently?

OpenStudy (anonymous):

this one is easier \(D=4J \) and \(D+10=20+2(J+10)\)

OpenStudy (anonymous):

but that's not one of the options, would it be c though bc that has the same ending just in a different order

OpenStudy (anonymous):

solve via \[4J+10=20+2(J+10)\] it is the same as C since \(2(J+10)+20=20+2(J+10)\)

OpenStudy (anonymous):

\[4J+10=20+2(J+10)\] \[4J+10=20+2J+20\] \[4J+10=2J+40\] \[2J=30\\ J=15\]

OpenStudy (anonymous):

son is 15, father is 60

OpenStudy (anonymous):

thank you so much for all of your help

OpenStudy (anonymous):

yw, good luck on your math

OpenStudy (anonymous):

thank you! good night (: